### v1

# The role of porosity in the transition to inertial regime in porous media flows

Abstract In this work we investigate the fundamental physical mechanism of the transition from Darcy to inertial (Darcy-Forchheimer) regime in the flow through porous media from the perspective of the vortex formation and growth. We focus on investigating the influence of porosity on the tortuosity–Reynolds number relation for periodic stochastic and simple cubic systems by numerically solving the Navier-Stokes equations in the pore-scale. We observe that the tortuosity defined by integrals over the fluid volume exhibits similar behavior in both types of systems. We also show that the tortuosity based on averaging of the streamlines length diverges from the volume-integrated one when the inertia onset takes place. We show that the discrepancy between those two tortuosities at increasing Reynolds number carries information about the geometrical confinement within the porous sample and is largely governed by the growth of the volumes occupied by vortices and their kinetic energy. Our results shed new light on the fundamental mechanisms of the transition between Darcy and the inertial regime and bring a better understanding of the role of the porous media geometry in this transition.
Keywords: porous media, inertial flows, Darcy-Forchheimer, tortuosity
1 Introduction
Many real-life flows through porous media occur at velocities at which the effects caused by fluid inertia cannot be omitted. Such scenarios are of much importance for gas extraction [1, 2], mixing and reactive processes [3], or flow through network-like porous media [4], e.g. the circulatory system [5, 6]. It is thus crucial from scientific and application point of view to understand well the physical principles of the transition of flow through porous media from the creeping flow (Darcy) regime to the inertial regime and the behavior of the flow within the latter. Of special interest are the relations between the pore-scale hydrodynamics and the effective (volume-averaged) hydrodynamic coefficients of a porous medium, which serve as a link connecting the fundamental fluid dynamical processes to the more practical quantities characterizing those systems.
In the Darcy regime, the Stokes equations are used to model the mass and momentum transport within the pore space. In order to obtain the average picture of the transport in this regime, upscaling [7, 8] is performed which results in the Darcy’s law [9]
$\boldsymbol{U}=\frac{k_{0}}{\nu}\left(-\langle\nabla P\rangle+\rho\boldsymbol{g}\right)$ (1)
stating the proportionality between the forcing term (mean pressure gradient $-\langle\nabla P\ \rangle$ and/or body force $\rho\boldsymbol{g}$) and the superficial flow velocity $\boldsymbol{U}$. The constant $k_{0}$ is the (intrinsic) permeability, characteristic for an isotropic porous system as $-\langle\nabla P\rangle$ and $\rho\boldsymbol{g}$ vanish, and $\nu$ is the fluid kinematic viscosity. In the inertial regime, the contribution of the convective derivative in the pore-scale momentum transport becomes non-negligible. Thus, the full Navier-Stokes equations need to be considered, yielding the Darcy-Forchheimer law [10]
$\boldsymbol{U}+\beta\rho\boldsymbol{U}|\boldsymbol{U}|=\frac{k_{0}}{\nu}\left(-\langle\nabla P\rangle+\rho\boldsymbol{g}\right)$ (2)
where $|\boldsymbol{U}|$ is the magnitude of the superficial velocity, $\beta$ is the free coefficient and the porous medium is again assumed to be isotropic. The transition from the Darcy to the inertial regime is often associated with the significant reorganization of the velocity and pressure fields within the pore spaces. Among many aspects of such reorganization, it was observed that the fluid kinetic energy becomes more evenly distributed in the pore space [11], which has to do with significant mass flow through an increasing number of channels and confinement of more kinetic energy within the vortices [12, 13, 14]. Our recent studies on high- [12] and low-porosity [14] porous media suggest that the porosity is one of the factors determining the dynamics of the transition from Darcy to inertial regime. Nevertheless, the exact mechanism of how the pores geometry impacts the onset of inertia is still unknown.
For the quantification of the inertial effects in porous media, a variety of measures are used. Fitting Eqs. (1) and (2) to the experimental data allows to asses the permeability and the inertial correction coefficient and to distinguish between various inertial subregimes [13, 15, 11]. The measurements of the hydrodynamic drag force exerted on the solid fraction of the porous medium can also be used as the inertia onset indicator [16, 17]. Another quantity, commonly used in the porous media studies is the tortuosity $T$, in certain cases related to the permeability through the Carman-Kozeny relation [18, 19] as $k_{0}\propto T^{-1}$. Although originally introduced as an effective parameter describing solely the porous medium geometry, the notion of tortuosity was later extended to encompass factors other than geometry [20], and even to non-hydrodynamic phenomena [21]. Due to this, in its nearly 100-year’s history, several definitions of the tortuosity has been proposed and the discussion on their range of applicability and relation to each other is still ongoing [22, 23, 24]. Majority of the hydrodynamic (hydraulic) tortuosity definitions define it as an average of the lengths of the streamlines obtained from the pore-scale velocity field. However, it was proven that one can calculate the tortuosity directly from the pore-scale velocity field, without the need to generate streamlines beforehand [25]. This observation proved to be useful in low-porosity systems where even at low superficial velocities recirculation zones are abundant and the automatic seeding of the streamlines poses a big challenge. However, this equivalence between the two approaches to calculate the tortuosity no longer holds in the inertial regime, where the impact of the recirculation zones on the pore-scale flow is much larger than in the Darcy regime [25]. This introduces even more ambiguity into the description and explanation of the inertia onset in porous media.
The aim of this work is to study how the interaction between the pores geometry and the pore-scale velocity drive the transition of porous media flow from the Darcy to the inertial regime. We investigate this transition quantitatively via two definitions of the tortuosity – one formulated using volume integrals of the velocity field, and the other, using the velocity streamlines. We explain why their dependence on the superficial flow velocity changes with the porosity of the porous sample. We show that those discrepancies, although seemingly contradictory to each other, provide information about the interaction between the geometry of the pores and the hydrodynamics in the pore-scale. In particular, we highlight that the main factor responsible for the behavior of the tortuosity during the Darcy-inertial transition is the dynamics of the growth of the recirculation zones in terms of their size and kinetic energy.
2 Methods
2.1 Physical and numerical setup
2.1.1 Governing equations and boundary conditions
We solve the steady-state incompressible Navier-Stokes equations in the volume of fluid $\Omega$ fully saturating the space between the solid phase of porous samples of side length $L$, with Dirichlet and Neumann boundary conditions on the solid walls $\partial\Omega$ for the velocity $\boldsymbol{V}$ and pressure $P$, respectively, and periodic boundary conditions along all three Cartesian coordinates
$\begin{array}[]{rll}\nabla\cdot\boldsymbol{V}=&\!\!\!\!0&\text{in }\Omega,\\ \boldsymbol{V}\>(\nabla\cdot\boldsymbol{V})=&\!\!\!\!-\nabla\dfrac{P}{\rho}+\boldsymbol{g}+\nu\nabla^{2}\boldsymbol{V}&\text{in }\Omega,\\ \boldsymbol{V}=&\!\!\!\!\boldsymbol{0}&\text{on }\partial\Omega,\\ \boldsymbol{n}\cdot\nabla P=&\!\!\!\!0&\text{on }\partial\Omega,\\ \boldsymbol{V}(\boldsymbol{x})=&\!\!\!\!\boldsymbol{V}(\boldsymbol{x}+\{0,\pm L\}^{3}),\\ P(\boldsymbol{x})=&\!\!\!\!P(\boldsymbol{x}+\{0,\pm L\}^{3}),\end{array}$ (3)
where the acceleration is aligned with the $x$-axis, $\boldsymbol{g}=[g,0,0]$ and $\boldsymbol{n}$ is the local unit normal vector on the solid boundary. We adopt the boldcase notation for the vectors and refer to the Cartesian coordinates with indices $0,1,2$, e.g. $\boldsymbol{V}\equiv[V_{0},V_{1},V_{2}]$. We denote the magnitude of a vector with a non-boldcase font, e.g. $|\boldsymbol{V}|\equiv V$.
It is useful to define at this point the notion of the percolating volume $\Omega_{p}\subset\Omega$. It is the volume consisting of all the streamlines that connect either two of the periodic boundaries of the fluid domain and are continuous within the porous sample (further called the percolating streamlines). The volume consisting of all fluid points which are not in $\Omega_{p}$ will be called the recirculation (or vortex) volume $\Omega_{v}$, i.e. $\Omega_{v}=\Omega\>\backslash\>\Omega_{p}$. Such definition of $\Omega_{v}$ is consistent with the definition of a steady-state vortex as the maximum set of nested closed streamlines used in previous works [26, 27]. With such division of $\Omega$, the boundary between the percolating and the recirculation volume, $\partial\Omega_{pv}$ can be clearly defined. A schematic representation of the percolating and recirculating volumes and the boundary between them is shown in Fig. 1.
Figure 1: A schematic representation of the percolating volume ($\Omega_{p}$, red), recirculation volume ($\Omega_{v}$, green) and the boundary between the two ($\partial\Omega_{pv}$, yellow) in a flow through a periodic porous sample (obstacles depicted gray). The curves with arrows along their length represent the percolating and the recirculating streamline. The dashed edges of the large square represent the periodic boundary.
2.1.2 Stochastic porous medium
The two stochastic porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We generate the geometry of the solid phase by dividing the periodic cell into $16^{3}$ cubes and subsequently marking each of them as solid with probability equal to the desired porosity $\phi$ of the sample. We consider values $\phi=0.7$ and $\phi=0.9$. The visualizations of the solid phase of the two samples are shown in Fig. 2.
To solve the governing equations in the stochastic samples we use Lattice Boltzmann Method [28, 29] implemented for the GPU execution using Template Numerical Library [30, 31]. We use D3Q19 velocity set for $\phi=0.7$ sample and D3Q27 for $\phi=0.9$ sample. Two-relaxation time (TRT) collision kernel [32] with the ‘magic parameter’ $\Lambda=1/4$ and the symmetric relaxation time $\tau^{+}=0.505$ is used, along with the lattice speed of sound $c_{s}=1/\sqrt{3}$. The discretization in space consists of $N=304$ lattice sites along each direction for $\phi=0.7$ sample and $N=160$ lattice sites along each direction for $\phi=0.9$ sample. The no-slip walls are implemented via the halfway-bounceback method [28, 29]. In the simulations with the highest Reynolds number, in the final timestep, the maximal value of either velocity component for either samples did not exceed $3.73\cdot 10^{-2}$ and the density variation $|1-\rho|$ did not exceed $1.73\cdot 10^{-3}$, all in lattice units.
The steady-state is assumed to be reached when the relative difference between the mass fluxes along $x_{0}$-direction per unit time measured at $x_{0}=0.25$ defined as
$\Delta Q_{0}=\frac{|Q_{0}^{t}-Q_{0}^{t-10^{4}\Delta t}|}{Q_{0}^{t}}\cdot\frac{1}{10^{4}\Delta t}$ (4)
and measured every $10^{4}$ timesteps falls below $10^{-2}$. In the above equation, $Q_{0}^{t}$ denotes the mass flux along $x_{0}$-direction at time $t$.
Figure 2: Visualization of the geometries of the solid phase of the two stochastic porous samples considered in this work. The porosities are $\phi=0.7$ (left) and $\phi=0.9$ (right). The black-edged cube denotes the boundaries of the periodic cell.
2.1.3 Simple cubic porous medium
The two simple cubic (SC) porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. The spherical obstacle is placed in the center of the periodic cell. We consider the obstacle radii $r=0.6526$ and $r=0.062$, resulting in porosities $\phi=0.1$ and $\phi=0.999$, respectively. For the convenience of visualization and meshing, the actual computational geometry for the lower-porosity sample considers the spherical obstacles placed in each of the periodic cell’s corners, i.e. the periodic cell is shifted by a vector $[L/2,L/2,L/2]$, The visualizations of the solid and fluid phases of the two samples are shown in Fig. 3.
We solve the governing equations with openFOAM [33]. We use the SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) algorithm [34] for the pressure-velocity coupling, implemented in simpleFOAM solver. The convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. The space discretization is generated by first creating the regular cubic grid with the discretization parameter $h=1/150$ using blockMesh and then adjusting it to the actual fluid volume geometry with snappyHexMesh.
Figure 3: Visualization of the geometries of the solid (gray) and fluid (blue, opaque) phases of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (left) and $\phi=0.999$ (right). The black-edged cube denotes the boundaries of the periodic cell. In the case of lowe-porosity sample only a few chosen obstacles residing in the periodic cell were chosen for clarity.
2.2 The used definitions of the tortuosity
To quantify the inertial effects in the flow, the present work considers two definitions of tortuosity. The first one, $T_{s}$, is defined as [23, 25, 22]
[…]
where $\tilde{\lambda}(\boldsymbol{x})$ is the normalized tortuosity of a streamline passing through a point $\boldsymbol{x}\in A\cap\Omega$
$\tilde{\lambda}(\boldsymbol{x})=\frac{\lambda(\boldsymbol{x})}{L_{\text{ref}}}$ (6)
with $L_{\text{ref}}$ denoting the reference length, unless otherwise stated we use $L_{\text{ref}}=L$. We will refer to this definition as the streamilne-based tortuosity. The integration in Eq. (5) is performed over the intersection of a plane $A$ perpendicular to the mean flow direction and the fluid volume $\Omega$. As noted by [25], the contribution of the recirculation zones to $T_{s}$ vanishes, thus Eq. (5) can be integrated over the intersection of the plane perpendicular to the mean flow direction and the percolating volume of the fluid
[…]
although in numerical evaluations of $T_{s}$ it is more convenient to place $A$ such that it intersects only the percolating volume. In this, and all further integrations, one can equivalently assume that $\boldsymbol{V}=\boldsymbol{0}$ inside the solid phase and include the obstacles in the integrations. $T_{s}$ integrates the tortuosities of streamlines seeded from $A$ with weighting proportional to the flux in the mean flow direction associated with each streamline. Such weighting considers the streamlines of large flux, and thus carrying more fluid mass in a unit time, as more important for the value of the tortuosity, which increases the impact of the preferential channels of the flow [11]. It was proven in [25] that when there are no recirculation zones in the flow, $T_{s}$ is equivalent to the other tortuosity considered in the present work
[…]
where the integration goes over the whole fluid volume $\Omega$. We will refer to this definition as the volume-integrated tortuosity. The notation $I_{\alpha}[\>\cdot\>]$ is a shorthand for the integral of some quantity over the fluid volume or its subset, i.e.
[…]
and will be used further in the text. When the recirculation zones are present in the flow, the two latter tortuosities are not equal, rather the inequality holds [25]
$T_{s}\leq T_{\Omega}.$ (10)
Out of the two discussed tortuosities, $T_{s},V_{0}$ and $T_{\Omega}$, the latter turns out to be significantly easier to calculate numerically. This is because of its locality – it does not make use of streamlines, which, due to numerical errors, can easily cross the boundary $\partial\Omega_{pv}$ between the percolating and recirculation volume during the integration along the fluid velocity field characteristics. This problem becomes especially pronounced in low-porosity porous media and at high Reynolds numbers, where a considerable portion of the fluid is occupied by the recirculation zones.
2.3 Post processing of the results
Generation of streamlines
[…]
The initial positions of the tracers depend on the task at hand. In the case of $\phi=0.1$ SC sample, for the calculation of the streamline-based tortuosity $T_{s}$, we place them uniformly at the fluid part of the $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers along $y$- and $z$-direction. To find the recirculation/percolating volumes, we place the tracers only in the vicinity of the solid boundaries at $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers. In both cases, streamlines that do not connect the inlet and the outlet, hit the solid wall, or have negative increment in $x_{0}$-direction are not included in the further analysis.
In case of $\phi=0.999$ SC sample, for the calculation of $T_{s}$ we place the tracers on a square grid of step $h_{s}=10^{-2}$ at the $x_{0}=0$ plane and dismiss the streamlines that do not fulfill the previously stated criteria. Due to numerical errors, the mean values of $x_{1}$- and $x_{2}$-velocities in this case are not identically zero, although they are two orders of magnitude lower than the mean $x_{0}$-velocity. It does not impact the calculation of $T_{\Omega}$ significantly, however causes the nearly-straight streamlines to experience a drift in the direction perpendicular to $x_{0}$. This results in their elongation form $L$ by a factor of magnitude comparable to the value of $T_{s}-1$. Due to this, for the higher-porosity SC sample, in Eq. (6) we use the distance between the streamline’s start and end as the normalization length, rather than the porous sample’s side length $L$. For the determination of the recirculation/percolating volume we use two sets of the initial positions of the tracers. To properly capture the flow separation on the rear part of the sphere’s surface we place the tracers upwind from the sphere, in the vicinity of the domain’s centerline $x_{1}=x_{2}=1$. After the integration, we apply the same criteria to dismiss certain streamlines as previously. To properly capture the saddle point at the end of the recirculation zone in the wake of the sphere, we place the tracers downwind from the sphere (also in the vicinity of the centerline) and integrate them in the upwind direction. We do not apply the previously described streamline exclusion criteria, rather we stop the integration before the streamlines reach the separation ring on the sphere’s surface, to avoid them falling into the recirculation zone due to the numerical errors. We use the separation $h_{s}=5\cdot 10^{-4}$ between the streamline starting positions along $y$- and $z$-direction for the identification of the recirculation/percolating volume.
Determining the recirculation and percolating volumes of fluid
To determine which points of the fluid domain $\Omega$ of the SC samples belong to the recirculation or percolating volume we divide the periodic cell into $N_{H}^{3}$ cubic cells and consider only those occupied by the fluid. We then mark those of the cells which contain at least one point of the streamlines generated for this purpose. Having obtained the cluster of the marked cells, we consider those lying on the cluster’s boundary (i.e. having at least one non-marked neighbor), near the expected position of the boundary between the recirculation and percolating volume, as the approximation of this boundary. Lastly, all the cells lying on either side of the approximate boundary are marked as belonging to the recirculation or the percolating volume. We use the values of $N_{H}=160$ for both systems.
Calculation of $T_{\Omega}$ constituents
To calculate the integrals of $V_{0}$ and $V$ over the whole volume of fluid, present in the definition of the volume-inegrated tortuosity $T_{\Omega}$, in the case of stochastic PM, the first-order method, i.e. summing the values of the hydrodynamic fields from each lattice point, is used. In the case of SC samples we use the ParaView utility IntegrateVariables [37, 38]. For the integration of $V_{0}$ and $V$ in the recirculation or the percolating volume we use the first order method, where the values of the fields are interpolated to centers of the indicator function using the same RBF-FD method as for the generation of streamlines.
3 Results
Difference in tortuosity–Reynolds number relation for porous samples of different porosities.
We first investigate the transition from the Darcy to inertial regime for two stochastic porous media (PM) samples of porosities $\phi=0.7$ and $\phi=0.9$. Fig. 4 shows the dependence of the tortuosity $T_{\Omega}$ defined in Eq. (8) on the Reynolds number
$Re=\frac{\sqrt{k_{0}}\langle V_{0}\rangle}{\nu\phi},\quad\langle V_{0}\rangle=\frac{I[V_{0}]}{I[1]}$ (11)
where $k_{0}$ is the permeability at $Re_{k}\rightarrow 0$. We note that for $Re_{k}$ values beyond the investigated range the solution went unsteady (transient). We observe that in the Stokes regime the tortuosity for each sample have a constant value, $1.07$ for $\phi=0.9$ and $1.175$ for $\phi=0.7$. As $Re$ increases, the tortuosity first decreases and then increases for both samples. However, the magnitude of the initial decrease and subsequent increase is considerably different for the two porosities. In the case of higher-porosity sample, those the decrease is about $0.02$ ($2/7$ of the initial value above unity) and the increase is about $0.01$. For the lower-porosity sample, these values are $0.003$ ($1/60$ of the initial value above unity) and $0.125$ ($0.7$ of the initial value above unity), respectively. Such differences between the minimal and maximal values achieved by $T_{\Omega}$ in the steady-state regime were previously observed in stochastic high-porosity media [12] and in low-porosity geological samples [14]. The onset of inertial effects is visible also in the visualizations of the velocity fields, as the weaker channeling at the outlets of the samples (Fig. 5) and the emergence of the recirculation zones in the wakes of the obstacles (Fig. 6).
Figure 4: $T_{\Omega}$ for stochastic porous media of porosity $\phi=0.7$ (left) and $\phi=0.9$ (right). The horizontal dashed lines denote the values of $T_{\Omega}$ in the limit $Re_{k}\rightarrow 0$. The inset in the left subplot shows the same data with different ranges of the axes.
Figure 5: Streamlines of the velocity field for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is along $x$ axis. The shown streamlines are generated using the streamTracer utility of ParaView.
Figure 6: Velocity vectors at a chosen sections for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is from left to right. For $\phi=0.7$, only every 3rd velocity vector in each direction is shown for clarity.
Splitting the volume-integrated tortuosity $T_{\Omega}$
In order to explain the difference between the tortuosity – Reynolds number relation described above, we first take a closer look at the definition of $T_{\Omega}$ and observe that both the numerator and denominator can be split into separate contributions coming from the percolating volume (denoted by subscript $p$) and the vortex, or recirculation, volume (denoted by subscript $v$)
$T_{\Omega}=\frac{I[V]}{I[V_{0}]}=\frac{I[V]_{p}+I[V]_{v}}{I[V_{0}]_{p}+I[V_{0}]_{v}}.$ (12)
According to the definition of the vortex from Sec. 2, the mean velocity of the vortex in steady-state flow is zero, i.e. $I[\boldsymbol{V}]_{v}=\boldsymbol{0}$, so the term $I[V_{0}]_{v}$ in the denominator can be omitted. Thus, we can rewrite Eq. (12) as
$T_{\Omega}=\frac{I[V]_{p}}{I[V_{0}]_{p}}+\frac{I[V]_{v}}{I[V_{0}]_{p}}.$ (13)
As the contribution of the recirculating zones to the value of the streamline-based tortuosity $T_{s}$ vanishes (cf. Eq. (7)), using the same procedure as in [25], it is possible to show that the integrals over the intersection of the surface $A$ and the fluid volume $\Omega$ in Eq. (7) can be transformed into the integrals over the percolating volume $\Omega_{p}$, i.e. $T_{s}=I[V]_{p}/I[V_{0}]_{p}$. This allows to rewrite Eq. (13) as
$T_{\Omega}=T_{s}+\frac{I[V]_{v}}{I[V_{0}]_{p}}.$ (14)
Such form of the volume-integrated tortuosity allows to clearly distinguish between the contributions from the percolating and the recirculation zones. The denominator, $I[V_{0}]_{p}$, is proportional to the Reynolds number (cf. Eq. (11)) and does not introduce any new unknown variable into our considerations.
As noted previously, the proper identification of the percolating and recirculating volumes in the fluid is very sensitive to stray streamlines (i.e. those that due to numerical errors cross the boundary between the percolating and recirculating volumes). During the present research we observed that even several stray streamlines can significantly alter the value of the streamline-based tortuosity. Because of this, we conduct the numerical investigation of the two contributions to $T_{\Omega}$ on a simple cubic (SC) systems and use thus obtained results to explain the behavior of $T_{\Omega}$ in the stochastic PM. SC samples can be thought of as models of particular fragments of a stochastic PM, at the same time having the geometry simple enough to allow for the accurate calculation of streamlines and identification of $\Omega_{p}$ and $\Omega_{v}$. To include in our analysis the cases with extremely high and low geometrical confinement, we consider SC systems of porosities $\phi=0.1$ (obstacle radius $0.6526$) and $\phi=0.999$ (obstacle radius $0.062$). Exemplary visualizations of the percolating volume in SC systems are shown in Fig. 8.
Figure 7: Visualizations of the percolating volumes in $\phi=0.1$ SC system for low (left) and high (right) $Re_{k}$. The gray surfaces are the obstacles’ walls and for the low-porosity system only a few chosen walls are shown for clarity. The visualizations use the histograms from which the percolating volumes were found. We show only the percolating volumes for clarity.
Figure 8: Visualizations of the recirculating volumes for $\phi=0.999$ SC system for intermediate (left) and high (right) $Re_{k}$ (for low $Re_{k}$ the recirculation zone is not present). The gray surfaces are the obstacles’ walls and for the low-porosity system only a few chosen walls are shown for clarity. The visualizations use the histograms from which the recirculation volumes were found. We show only the recirculation volumes for clarity.
Tortuosity as the inertia onset indicator in simple cubic (SC) systems
We begin our analysis by looking at the Reynolds number dependence of $T_{\Omega}$ and $T_{s}$ for SC sample, shown in Fig. 9. In the Darcy regime ($Re_{k}\rightarrow 0$), the values of both tortuosities stay constant and are compliant with each other, as shown in [25]. When the Reynolds number increases, the streamline-based tortuosity, $T_{s}$, decreases for both systems, which was observed in stochastic and periodic porous media by other authors [20, 13]. On the other hand, when the volume-based tortuosity, $T_{\Omega}$, is concerned, the inertial effects manifests themselves differently depending on the porosity, similarly to the stochastic porous media (cf. Fig. 4). In the lower-porosity case, its value decreases slightly in the range $Re_{k}\in[10^{-3};10^{-2}]$ to later increase significantly until the steady-state regime ends. Such behavior resembles the one observed for the lower-porosity stochastic porous media in Fig. 4 and also for a real-life porous sample [14]. For higher-porosity sample, the volume-based tortuosity follows the streamline-based one closely, to diverge from it only at $Re_{k}\approx 10^{-3}$. Similarly to the lower-porosity system, the initial decline of $T_{s}$ is followed by its rise, however to a much smaller extent. This is in line with the behavior of $T_{\Omega}$ for the higher-porosity stochastic porous media shown in Fig. 4, as well as a similar system investigated previously [12].
The strictly decreasing behavior of the streamline-based tortuosity can be attributed to the rectification of streamlines which follow the growth of the recirculation zones volumes in the cavities and obstacles wakes. The volume-integrated tortuosity exhibits a more complex behavior, dependent on the sample’s porosity. Eq. (14) shows that the difference between $T_{\Omega}$ and $T_{s}$ is related to the integral of velocity magnitude inside the recirculation zones. Thus, we will further investigate the fall and the rise of the volume-based tortuosity in terms of the volume of the recirculation zones and the field of the velocity magnitude therein.
Figure 9: $T_{\Omega}$ and $T_{s}$ for $\phi=0.1$ and $\phi=0.999$ SC samples. The inset in the left subplot shows the magnification of the $\phi=0.1$ plot to the range of $T_{s}$ values. The horizontal dashed lines represent the values of $T_{\Omega}$ calculated for vanishingly small $Re_{k}$.
Origin of the $T_{\Omega,v}$ decrease and rise – volume of the recirculation zones
Let us first focus on the numerator of the percolation-volume related term from Eq. (14), $I[V]_{v}/I[V_{0}]_{p}$. For convenience, we divide it by the total fluid volume, $I[1]$, i.e.
$\frac{I[V]_{v}}{I[1]}=\frac{I[1]_{v}}{I[1]}\cdot\frac{I[V]_{v}}{I[1]_{v}}$ (15)
where the first term on the right-hand side represents the vortex volume normalized by the total volume of fluid and the other one – the total velocity magnitude per unit recirculation zone volume. We consider each of those two components separately. Fig. 8 shows the visualization of the percolating volumes and recirculation zones for low- and high-porosity SC samples, respectively, for two different Reynolds numbers each. In both systems, as $Re_{k}$ rises, the recirculation zones grow in volume, which in the case of low-porosity system is visible through the percolating volume shrinking to a straight jet connecting the inlet and the outlet. To quantify the changes in the recirculation zones size, we plot their volume, $I[1]_{v}$, normalized by the total volume of fluid, $I[1]$, in Fig. 10. In the low-porosity system, there already exist recirculation zones in the Darcy regime. The increase in their volume starts at $Re_{k}\approx 10^{-3}$, which corresponds to the start of $T_{\Omega}$ decrease in Fig. 9. The vortex volume then continues to grow until the end of the steady state regime, with significant slowdown for $Re_{k}>4\cdot 10^{-2}$, the value at which the slope of the $T_{\Omega}$ rise in Fig. 9 decreases. This reduction in the recirculation volume growth pace indicates the strong geometrical confinement for the vortex development. The total increase in the volume of the recirculation zone is about 46% of the fluid volume, or – equivalently – $46/90\approx 0.51$ of the volume of the solid obstacles. On the other hand, in the high-porosity system there is no recirculation in the Darcy regime. The increase of its volume starts at $Re_{k}\approx 2\cdot 10^{2}$, which is about two decades later than the start of $T_{\Omega}$ decrease for this system in Fig. 9. As the recirculation volume starts to grow, it does so with a slightly increasing slope until the end of the steady-state regime. The increase in the volume of the recirculation zone in this case is about $0.63$ of the solid obstacle volume. The growth of the recirculation zones in the case of high-porosity SC systems was highlighted by other authors [17] as the reason for lower drag compared to the flow over a single obstacle.
Figure 10: Relative vortex volumes, $I[1]_{v}/I[1]$, for the two studied SC systems.
Origin of the $T_{\Omega,v}$ decrease and rise – energy confinement in the recirculation oznes
The increase of the recirculation zone volume can explain the decrease of $T_{\Omega}$ because it directly implies the rectification of all the percolating streamlines and thus the decrease of $T_{s}$ in Eq. 14. However, it fails to address the rise of the volume-based tortuosity at the highest Reynolds numbers. This is why now we turn our attention to the integral of the velocity magnitude inside the vortex per vortex volume, $I[V]_{v}/I[1]_{v}$ (further in the text we shall refer to this quantity as the specific velocity magnitude in vortex, for brevity). Its dependence on the Reynolds number is shown in Fig. 11 for both SC systems. We note that since in Darcy regime there was no recirculation zone present in the higher-porosity SC system, we assumed the value of the specific velocity magnitude in vortex to be zero in those cases. For both SC systems, the specific velocity magnitude in vortex increases with Reynolds number. In the low-porosity sample, the rise proceeds from the lowest $Re_{k}$ with the first order slope ($p=1$) and experiences a jump in the slope around $Re_{k}=10^{-2}$. It later attains the slope of the order approximately one again, in the range of $Re_{k}$ corresponding to those where the growth of $T_{\Omega}$ slows down (cf. Fig. 9). The initial first-order dependence is explained by the linearity of the Stokes equation and proportionality between the body force magnitude $g$ and the mean flow velocity $\langle\boldsymbol{V}\rangle$ in the Darcy regime (i.e., when $g$ is increased, the velocity field inside the vortex is only scaled proportionally by a constant). The subsequently higher slope indicates that a non-trivial mechanism involving the momentum transfer to the recirculation zone and the growth of its volume comes into play in the inertial regime. At the same time, an analogy between the Darcy regime and the inertial regime above $Re_{k}=3\cdot 10^{-2}$ can be drawn in a sense that both are characterized by the recirculation zone growing more in kinetic energy, rather than volume. This further highlights the importance of the geometrical confinement in the emergence of inertial effects in low-porosity samples [14]. In the high-porosity system, the start of the rapid growth of the specific velocity magnitude in vortex happens roughly at $Re_{k}$ corresponding to the start of the growth of the recirculation zone volume (cf. Fig. 10). It is interesting that even though the integral of the velocity magnitude over the whole fluid volume, $I[V]$, achieved at the highest Reynolds number is about $36$ times larger in the high-porosity system, the specific velocity magnitude in vortex is over 2 times higher in low-porosity system ($\approx 140$ vs. $\approx 60$). This suggests that the strong geometrical confinement leads to the entrapment of relatively large amounts of the kinetic energy in the recirculation zones, quantified also by other authors using the variance of the kinetic energy [11, 13, 12, 14].
In the end, in the low-porosity stochastic and SC systems, the greater kinetic energy confinement in the recirculation zones and the cease of their volume growth at high Reynolds numbers lead to the more pronounced increase of $T_{\Omega}$, compared to the higher-porosity counterparts.
Figure 11: Integral of the absolute value of velocity in the vortex per vortex volume for SC systems (for brevity called the specific velocity magnitude in vortex in the text, $I[V]_{v}/I[1]_{v}$). The inset in the left subplot shows the same data but on a log-log scale.
4 Conclusions
The presented findings on the differences in behavior of the streamline-based and volume-integrated tortuosity during the Darcy-inertial transition advance the understanding of the mechanisms behind the inertia onset in the studied porous systems. We show that porosity and the associated geometrical confinement are two of the key factors determining the behavior of the system in the inertial regime. The growth of the recirculation zones volume and the corresponding rectification of the percolating streamlines are captured by both types of tortuosity via their decrease in the initial stage of inertia onset. The later dynamics of the inertial effects emergence is strongly dependent on the system’s geometry. In low-porosity media, the geometrical confinement prevents the recirculation zones from growing freely in volume, thus mainly the kinetic energy confined within them increases. This is represented by the saturation of the streamline-based tortuosity at its minimal value, while the volume-integrated tortuosity grows rapidly in this regime. This is because the volume-integrated tortuosity contains information about the whole fluid volume, while the streamline-based one – only about the percolating volume. On the other hand, in the high-porosity systems the growth of the recirculation zones volume is not limited by the solid walls. Due to this, the increase of the volume-integrated tortuosity is less pronounced compared to the low-porosity samples, and the range of Reynolds numbers over which the two definitions of tortuosity are compliant with one another can extend almost to the end of the steady-state flow regime.
The investigated physical phenomena are closely related to the tools used to quantify them. The presented results show the utility of the two discussed measures of the inertial effects in porous media – streamline-based and volume-integrated tortuosity. Each of them separately carries information about various forms of the inertial effects in the flow – $T_{s}$ is more sensitive to the changes occurring solely in the percolating volume, while in $T_{\Omega}$ the recirculation zones contribution is also visible. The difference between the values of the two, as well as the magnitude of the growth of $T_{\Omega}$, can be treated as an indicator of the confinement of the recirculation zones. The presented results also highlight which aspects of the inertia onset in porous media are better captured by which definition of tortuosity and why, which will allow for a more sensitive choice between the two in more application-oriented problems. One may stipulate, for instance, that the information about the recirculation zones carried by the volume-integrated tortuosity might prove useful when the mixing and reaction phenomena in the porous sample are of interest. On the other hand, the exclusive contribution of the percolating volume to the streamline-based tortuosity may carry more valuable information for studies concerned with the mass transport through the sample, e.g. during gas extraction, where the gaseous phase is often transported through preferential channels and the residuals of the fluid phase occupy the dead-end pores and cavities.
5 Acknowledgments
Funded by National Science Centre, Poland under the OPUS call in the Weave programme 2021/43/I/ST3/00228. This research was funded in whole or in part by National Science Centre (2021/43/I/ST3/00228). For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. D. Strzelczyk and M. Matyka acknowledge the financial support from the Slovenian Research And Innovation Agency (ARIS) research core funding No. P2-0095.
Appendix A Relation to other measures for the onset of inertia
To give a better perspective on our present findings, for the discussed stochastic and simple cubic systems we calculate other the measures of the inertia onset. Fig. 12 shows the relation between the mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term. A comparison with Figs. 4 and  9 shows that for the higher-porosity samples of each type the Reynolds number at which the mean velocity–forcing term relation becomes nonlinear corresponds roughly to the beginning of the fall of the tortuosities ($Re_{k}\approx 0.52$ and $5.75$ for the stochastic and simple cubic system, respectively). For the lower-porosity samples, the transition to the nonlinear dependence occurs later, around the minimum of $T_{\Omega}$ ($Re_{k}\approx 0.56$ and $9.38\cdot 10^{-3}$, for the stochastic and simple cubic system, respectively). Note also that for the low-porosity simple cubic sample, the divergence from linearity in $\langle V_{0}\rangle$$g$ relation is visible the least from among all four samples. This is in contrast to the $T_{\Omega}$, which for the same sample experiences the biggest jump in value between the Darcy regime and the highest $Re_{k}$. This suggests that the volume-integrated tortuosity might be a good indicator for the onset of inertia for extremely low-porosity systems.
Figure 12: Mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first-order scaling slopes anchored at the data points with the lowest value of $g$ (for some systems those points are not visible at the presented axes ranges). The values at the arrows show the $Re_{k}$ for the data points around which the relation becomes non-linear. The acceleration and velocity for stochastic systems are given in lattice units.
Fig 13 shows the dependence of the inertial correction $f_{c,0}$ defined as [15]
$f_{c,0}=-\left(\frac{\langle\boldsymbol{V}\rangle}{|\langle\boldsymbol{V}\rangle|}-\frac{k_{0}\boldsymbol{g}}{\mu|\langle\boldsymbol{V}\rangle|}\right)\cdot\hat{\boldsymbol{x}}_{0}=-1+\frac{k_{0}g}{\mu\langle V_{0}\rangle}$ (16)
on the Reynolds number $Re_{k}$ for the studied stochastic and simple cubic systems. In Eq. (16), $\hat{\boldsymbol{x}}_{0}$ is the unit vector aligned with the $x_{0}$-axis and the last equality holds under the assumption of the mean flow aligned with the $x_{0}$ direction, namely $\langle\boldsymbol{V}\rangle=\langle V_{0}\rangle\hat{\boldsymbol{x}}_{0}$. The weak inertia regime is characterized by the $f_{c,0}\propto Re_{k}^{2}$ scaling, while the strong inertia regime – by $f_{c,0}\propto Re_{k}$ scaling. In the case of stochastic samples, both regimes are vaguely visible, which suggest that the transition between the two regimes takes place over the majority of the considered Reynolds numbers. In the case of simple cubic systems, the weak inertia regime is clearly visible for both porosities. For the lower-porosity sample, in the lowest $Re_{k}$ cases the scaling becomes below-second order, which was also observed by other authors [15]. The deviation from the weak inertia regime apperas approximately at $Re_{k}=10^{-2}$, for which the $T_{\Omega}$ reaches its minimum (cf. Fig 9). For the highest Reynolds numbers, the relation changes into sublinear, however it endures over a short range of $Re_{k}$ before the steady-state regime ends. It is possible that one observes the regime above strong inertia where $f_{c,0}\propto Re_{k}^{2}+Re_{k}$. One might argue though that this regime is related to the reduction in the increase of the slope of the volume-based tortuosity at the highest $Re_{k}$. For the higher-porosity simple cubic sample, relation becomes sub-quadratic around $Re_{k}=1$, the value at which the tortuosities in Fig. 9 begin to decrease. Also in this case, for the highest Reynolds numbers, the $f_{c,0}(Re_{k})$ relation is clearly non-linear. However, the highest-Reynolds number results should be regarded with caution as one operates close to the limit of the stability of the steady-state solver.
Figure 13: The inertial corection $f_{c,0}$ (Eq. (16)) versus the Reynolds number for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first- and second-order scaling slopes for the stochastic samples and $1/2$– and second order scaling slopes for the simple cubic samples. The inset in the right subplot shows the magnification of the highest-$Re_{k}$ region for the lower-porosity sample.