### v2

# The role of porosity in the transition to inertial regime in porous media flows

Abstract In this work, we investigate the fundamental physical mechanism of the transition from Darcy to inertial (Darcy-Forchheimer) regime in steady-state flows through porous media, with the focus on vortex formation. We investigate their influence on the tortuosity–Reynolds number relation during this transition for systems of various porosities. We do so by numerically solving the Navier-Stokes equations within the pore-scale of simple cubic systems and relating the observations made therein to stochastic systems of more complex geometry. We observe that the tortuosity defined by integrals over the whole fluid volume behaves similarly in both types of systems. At the same time, in simple cubic systems, the tortuosity based on averaging of the length of the streamlines diverges from the volume-integrated one when the inertia onset takes place. We show that the discrepancy between those two tortuosities at increasing Reynolds number carries information about the dynamics of the vortex growth in the system. We stipulate that those dynamics are directly governed by the porosity. Our results highlight the utility of various definitions of tortuosity as measures of inertia in porous media flows and explain the reasons for the differences between those definitions. This can lead to a more sensible choice of inertia indicators in more application-oriented problems.
Keywords: porous media, inertial flows, Darcy-Forchheimer, tortuosity
1 Introduction
Many real-life flows through porous media occur at velocities at which the effects caused by fluid inertia cannot be omitted. Such scenarios are of much importance for gas extraction [1, 2], mixing and reactive processes [3, 4], or flow through network-like porous media [5], e.g. the circulatory system [6, 7]. It is thus crucial from the scientific and application point of view to well well understand the physical principles of the transition of flow through porous media from the creeping flow (Darcy) regime to the inertial regime and the behavior of the flow within the latter. Of special interest are the relations between the pore-scale hydrodynamics and the effective (volume-averaged) hydrodynamic coefficients of a porous medium, which serve as a link connecting the fundamental fluid dynamical processes to the more practical quantities characterizing those systems.
In the Darcy regime, the Stokes equations are used to model the mass and momentum transport within the pore space. In order to obtain the average picture of the transport in this regime, upscaling [8, 9] is performed which results in the Darcy’s law [10]
$\boldsymbol{U}=\frac{k_{0}}{\nu}\left(-\langle\nabla P\rangle+\rho\boldsymbol{g}\right)$ (1)
stating the proportionality between the forcing term (mean pressure gradient $\langle\nabla P\ \rangle$ and/or body force $\rho\boldsymbol{g}$) and the superficial flow velocity $\boldsymbol{U}$. The constant $k_{0}$ is the (intrinsic) permeability characteristic of an isotropic porous system as $\langle\nabla P\rangle+\rho\boldsymbol{g}$ term vanishes, and $\nu$ is the fluid kinematic viscosity. In the inertial regime, the contribution of the convective derivative in the pore-scale momentum transport becomes non-negligible. Thus, the full Navier-Stokes equations need to be considered, yielding the Darcy-Forchheimer law [11, 12]
$\boldsymbol{U}+\beta\rho\boldsymbol{U}|\boldsymbol{U}|=\frac{k_{0}}{\nu}\left(-\langle\nabla P\rangle+\rho\boldsymbol{g}\right)$ (2)
where $|\boldsymbol{U}|$ is the magnitude of the superficial velocity, $\beta$ is the free coefficient and the porous medium is again assumed to be isotropic. The transition from the Darcy to the inertial regime is often associated with the significant reorganization of the velocity and pressure fields within the pore spaces. Among many aspects of such reorganization, it was observed that the fluid kinetic energy becomes more evenly distributed in the pore space [13], which has to do with significant mass flow through an increasing number of channels and confinement of more kinetic energy within the vortices [14, 15, 16]. Our recent studies on high- [14] and low-porosity [16] porous media suggest that the porosity is one of the factors determining the dynamics of the transition from Darcy to inertial regime. Nevertheless, the exact mechanism of how the geometry of the pores impacts the onset of inertia is still unknown.
For the quantification of the inertial effects in porous media, a variety of measures are used. Fitting Eqs. (1) and (2) to the experimental data allows one to assess the permeability and the inertial correction coefficient and to distinguish between various inertial subregimes [15, 17, 13]. The measurements of the hydrodynamic drag force exerted on the solid fraction of the porous medium can also be used as the inertia onset indicator [18, 19]. Another quantity, commonly used in porous media studies, is the tortuosity $T$, in certain cases related to the permeability through the Carman-Kozeny relation [20, 21] as $k_{0}\propto T^{-1}$. Although originally introduced as an effective parameter describing solely the porous medium geometry, the notion of tortuosity was later extended to encompass factors other than geometry [22], and even to non-hydrodynamic phenomena [23]. Due to this, in its nearly 100-year history, several definitions of the tortuosity have been proposed, and the discussion on their range of applicability and relation to each other is still ongoing [24, 25, 26]. The majority of the hydrodynamic (hydraulic) tortuosity definitions define it as an average of the lengths of the streamlines obtained from the pore-scale velocity field. However, it was proven that one can calculate the tortuosity directly from the pore-scale velocity field, without the need to generate streamlines beforehand [27]. It becomes especially useful in low-porosity systems where even at low superficial velocities, recirculation zones are abundant and the automatic seeding of the streamlines poses a big challenge. However, this equivalence between the two approaches to calculate the tortuosity no longer holds in the inertial regime, where the impact of the recirculation zones on the pore-scale flow is much larger than in the Darcy regime [27]. This introduces even more ambiguity into the description and explanation of the onset of inertia in porous media.
The aim of this work is to study the behavior of tortuosity during the transition from the Darcy to the inertial regime. We do so to explain how the interplay between the pores’ geometry and the pore-scale velocity drives this transition in porous media flows. To describe the transition quantitatively, we use two definitions of the tortuosity – one formulated using volume integrals of the velocity field, and the other, using the velocity streamlines. We explain why their dependence on Reynolds number changes with the porosity of the porous sample. We show that those differences provide information about the interaction between the geometry of the pores and the hydrodynamics in the pore-scale. In particular, we highlight that the main factor affecting the tortuosity during the Darcy-inertial transition is the growth of the recirculation zones in size and kinetic energy.
2 Methods
2.1 Physical and numerical setup
2.1.1 Governing equations and boundary conditions
We solve the steady-state mass conservation (Eq. (3a)) and Navier-Stokes (Eq. (3b)) equations for an incompressible fluid in the pore-scale of three-dimensional porous samples. We consider cubic samples of side length $L$. The fluid fully saturates the pore space $\Omega$. On the solid walls, $\partial\Omega_{\text{solid}}$, we impose Dirichlet boundary conditions for the velocity $\boldsymbol{V}$ (Eq. (3c)) and Neumann boundary conditions for the pressure $P$ (Eq. (3d)). We also impose periodic boundary conditions along all three Cartesian coordinates (Eqs. (3e) and (3f))
$\nabla\cdot\boldsymbol{V}$ $=0$ $\quad\text{in }\Omega,$ (3a)
$\boldsymbol{V}\>(\nabla\cdot\boldsymbol{V})$ $=-\nabla\dfrac{P}{\rho}+\boldsymbol{g}+\nu\nabla^{2}\boldsymbol{V}$ $\text{in }\Omega,$ (3b)
$\boldsymbol{V}$ $=\boldsymbol{0}$ $\text{on }\partial\Omega_{\text{solid}},$ (3c)
$\boldsymbol{n}\cdot\nabla P$ $=0$ $\text{on }\partial\Omega_{\text{solid}},$ (3d)
$\boldsymbol{V}(\boldsymbol{x})$ $=\boldsymbol{V}(\boldsymbol{x}+\{0,\pm L\}^{3})$ $\text{in }\Omega,$ (3e)
$P(\boldsymbol{x})$ $=P(\boldsymbol{x}+\{0,\pm L\}^{3})$ $\text{in }\Omega.$ (3f)
The acceleration is aligned with the $x$-axis, $\boldsymbol{g}=[g,0,0]$, and $\boldsymbol{n}$ is the local unit normal vector on the solid boundary. We adopt the boldcase notation for the vectors and refer to the Cartesian coordinates with indices $0,1,2$, e.g. $\boldsymbol{V}\equiv[V_{0},V_{1},V_{2}]$. We denote the magnitude of a vector with a non-boldcase font, e.g. $|\boldsymbol{V}|\equiv V$.
It is useful to define at this point the notion of the percolating volume $\Omega_{p}\subset\Omega$. It is the volume consisting of all the streamlines that connect any pair of the periodic boundaries of the fluid domain and are continuous within the porous sample (further called the percolating streamlines). The volume consisting of all fluid points which are not in $\Omega_{p}$ will be called the recirculation (or vortex) volume $\Omega_{v}$, i.e. $\Omega_{v}=\Omega\>\backslash\>\Omega_{p}$. Such a definition of $\Omega_{v}$ is consistent with the definition of a steady-state vortex as the maximum set of nested closed streamlines used in previous works [28, 29]. With such division of $\Omega$, the boundary between the percolating and the recirculation volume, $\partial\Omega_{pv}$ can be clearly defined. A schematic representation of the volumes and the boundaries considered in this work is shown in Fig. 1.
Figure 1: A schematic representation of the percolating volume ($\Omega_{p}$, red), recirculation volume ($\Omega_{v}$, green) and the boundary between the two ($\partial\Omega_{pv}$, yellow) in a flow through a periodic porous sample (obstacles depicted striped gray with their boundary $\partial\Omega_{\text{solid}}$ plotted with solid black line). The curves with arrows along their length represent the percolating and the recirculating streamlines. The dashed edges of the large square represent the periodic boundary
2.1.2 Stochastic porous medium
The two stochastic porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We generate the geometry of the solid phase by dividing the periodic cell into $16^{3}$ cubes and subsequently marking each of them as fluid with probability equal to the desired porosity $\phi$ of the sample. We consider values $\phi=0.7$ and $\phi=0.9$ and obtain the actual porosities of $0.696$ and $0.897$, respectively. The visualizations of the solid phase of the two samples are shown in Fig. 2.
To solve the governing equations in the stochastic samples, we use Lattice Boltzmann Method (LBM) [30, 31] implemented for the GPU execution using Template Numerical Library [32, 33]. Due to the uniform discretization of space, LBM allows for simple treatment of complex boundaries and has been successfully used to model flows in porous media [34, 35, 36], We use D3Q19 velocity set for $\phi=0.7$ sample and D3Q27 set for $\phi=0.9$ sample, for better stability at higher velocities. Two-relaxation time (TRT) collision kernel [37] with $\Lambda=1/4$ and the symmetric relaxation time $\tau^{+}=0.505$ is used, along with the lattice speed of sound $c_{s}=1/\sqrt{3}$. The discretization in space consists of $N=304$ lattice sites along each direction for $\phi=0.7$ sample and $N=160$ lattice sites along each direction for $\phi=0.9$ sample. The no-slip walls are implemented via the halfway-bounceback method [30, 31]. In the simulations with the highest Reynolds number, in the final timestep, the maximal value of either velocity component for either sample did not exceed $3.73\cdot 10^{-2}$ and the density variation $|1-\rho|$ did not exceed $1.73\cdot 10^{-3}$, all in lattice units.
The steady-state is assumed to be reached when the relative difference between the mass fluxes along $x_{0}$-direction per unit time measured at $x_{0}=0.25$ defined as
$\Delta Q_{0}=\frac{|Q_{0}^{t}-Q_{0}^{t-10^{4}\Delta t}|}{Q_{0}^{t}}\cdot\frac{1}{10^{4}\Delta t}$ (4)
and measured every $10^{4}$ timesteps falls below $10^{-2}$. In the above equation, $Q_{0}^{t}$ denotes the mass flux along $x_{0}$-direction at time $t$.
Figure 2: Visualization of the geometries of the solid phase of the two stochastic porous samples considered in this work. The porosities are $\phi=0.7$ (left) and $\phi=0.9$ (right). The black-edged cube denotes the boundaries of the periodic cell.
2.1.3 Simple cubic porous medium
The two simple cubic (SC) porous samples considered in this work are confined within a cubic periodic cell of side length $L=1$. We consider porosities $\phi=0.1$ and $\phi=0.999$ corresponding to the obstacle radii $r=0.6526$ and $r=0.062$, respectively. For the convenience of visualization and meshing, the computational geometry for the lower-porosity sample considers the spherical obstacles placed in each of the periodic cell’s corners, and for the higher-porosity sample, we place a single obstacle in the center of the domain (see Fig. 3).
For SC samples, we solve the governing equations with openFOAM [38]. We use the SIMPLE (Semi-Implicit Method for Pressure-Linked Equations) algorithm [39] for the pressure-velocity coupling, implemented in simpleFOAM solver. In contrast to LBM used with the stochastic samples, SIMPLE allows one to solve directly for the steady-state solution, shortening the calculation time. For $\phi=0.1$, the convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. All considered systems reach this convergence criterion, and the verification of the steady-state solution for the system of the highest flow rate is performed with pimpleFOAM. The space discretization is generated by first creating the regular cubic grid with the discretization parameter $h=1/195$ using blockMesh and then adjusting it to the actual fluid volume geometry with snappyHexMesh (see Fig. 4).
For $\phi=0.999$, the convergence criterion based on residuals of the velocity and pressure is set to the values of $10^{-5}$ and $10^{-4}$, respectively. Three systems of the highest flow rate did not reach this criterion; however, the transient simulations performed with pimpleFOAM allow them to reach the steady state, and those data are used in the subsequent analysis. The determination of the steady state in those cases is based on the relative difference between the current and the estimated final value of the $x_{0}$-component of velocity in the bulk and in the wake of the obstacle. In the final iteration, this difference does not exceed $10^{-3}$. The final values of the velocities were estimated from the fit of the exponential function to the values of those velocities in time. We note that the $\phi=0.999$ system becomes unsteady in the range of $Re_{k}\in\>(5\cdot 10^{3};10^{4}]$. The discretization for $\phi=0.999$ systems solved using simpleFOAM is created the same way as for the $\phi=0.1$ systems. For the three highest flow rates solved with pimpleFOAM, we first create a regular cubic grid with the discretization parameter $h_{\text{bulk}}=1/75$ using blockMesh and then adjust it to the actual fluid volume geometry with snappyHexMesh with local refinement ($h_{\text{min}}=1/150$) near the sphere and in its wake (see Fig. 4). For all transient simulations (for both porosities), we choose the timestep $\Delta t$ such that the maximal Courant number in the system is below unity.
Figure 3: Visualization of the geometries of the solid (gray) and fluid (blue, opaque) phases of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (left) and $\phi=0.999$ (right). The black-edged cube denotes the boundaries of the periodic cell. In the case of the low-porosity sample, only a few chosen obstacles residing in the periodic cell are shown for clarity.
Figure 4: Visualization of the meshes of the two simple cubic porous samples considered in this work. The porosities are $\phi=0.1$ (two top subplots) and $\phi=0.999$ (two bottom subplots, for the uniform and refined discretization). The top-right subplot shows the magnified view of the top-left mesh. The black-edged cube denotes the boundaries of the periodic cell. Only $z<0.5$ half of each mesh is shown.
2.2 The used definitions of the tortuosity
To quantify the inertial effects in the flow, let us introduce three definitions of tortuosity, two of which will be used further in this work. They mainly make use of the streamlines, defined as curves $s$ tangent to the velocity field vectors at each of their points. Equivalently, for the steady state systems, a streamline can be defined as a path along which a massless particle is advected in the velocity field (see Fig. 5). Since the velocity field has only one value in each fluid point $\boldsymbol{x}\in\Omega$, the streamlines cannot cross each other, and thus, there exists a 1-to-1 mapping between the fluid points and the streamlines. In other words, a streamline can be identified uniquely by any of its points: $s\equiv s(\boldsymbol{x})$ for any $\boldsymbol{x}\in s(\boldsymbol{x})$. We will use this fact when introducing the definitions of tortuosities in the following part of this section.
Figure 5: Schematic representation of a streamline $s$ as a curve tangent to the fluid velocity field vectors (black arrows) in each of the streamline’s points (two chosen are shown here as black dots, $\boldsymbol{x}$ and $\boldsymbol{x^{\prime}}$). An infinitesimal increment along a streamline, $d\boldsymbol{s}$ (small blue arrow), is parallel to the local velocity vector, here normalized for convenience. Any of the streamline’s points identifies it uniquely ($s(\boldsymbol{x})\equiv s(\boldsymbol{x^{\prime}})$).
The first presented definition of the tortuosity, $T_{s}^{*}$, was originally defined as a weighted average of the tortuosities of all the streamlines in the system [25, 27, 24]
$T_{s}^{*}=\frac{\displaystyle\int\limits_{A\>\cap\>\Omega}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})\tilde{\lambda}(\boldsymbol{x})}{\displaystyle\int\limits_{A\>\cap\>\Omega}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})}.$ (5)
where $\tilde{\lambda}(\boldsymbol{x})$ is the tortuosity of a streamline passing through a point $\boldsymbol{x}\in A\cap\Omega$
$\tilde{\lambda}(\boldsymbol{x})=\dfrac{\lambda(\boldsymbol{x})}{L_{\text{ref}}},\quad\lambda(\boldsymbol{x})\equiv\!\!\!\!\!\!\!\!\int\limits_{\text{along }s(\boldsymbol{x})}\!\!\!\!ds$ (6)
with $\lambda(\boldsymbol{x})$ denoting the (curvilinear) length of the streamline, $ds$ the length of the infinitesimal increment along the streamline, and $L_{\text{ref}}$ denoting the reference length. Unless otherwise stated, we take the reference length to be equal to the length of the porous sample’s side, i.e., $L_{\text{ref}}=L$. The integration in Eq. (5) is performed over the intersection of a plane $A$ perpendicular to the mean flow direction and the fluid volume $\Omega$ (see Fig. 6, leftmost plot). Also, $A$ must be intersected by all the streamlines in the system. As it is ambiguous to tell the length of a looping streamline inside a recirculation zone, the integration in Eq. (5) is usually limited only to the intersection of the plane perpendicular to the mean flow direction and the percolating volume of the fluid (see Fig. 6, middle plot)
$T_{s}=\frac{\displaystyle\int\limits_{A\>\cap\>\Omega_{p}}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})\tilde{\lambda}(\boldsymbol{x})}{\displaystyle\int\limits_{A\>\cap\>\Omega_{p}}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})},$ (7)
although in numerical evaluations of $T_{s}$ it is more convenient to place $A$ such that it intersects only the percolating volume. In the present work, we will use the definition from Eq. (7) and will refer to it as the streamilne-based tortuosity. In this, and all further integrations, one can equivalently assume that $\boldsymbol{V}=\boldsymbol{0}$ inside the solid phase and include the obstacles in the integrations. $T_{s}$ integrates the tortuosities of the percolating streamlines seeded from $A\cap\Omega_{p}$ with weighting proportional to the flux in the mean flow direction associated with each streamline. Such weighting considers the streamlines of large flux, and thus carrying more fluid mass in a unit time, as more important for the value of the tortuosity, which increases the impact of the preferential channels of the flow [13]. It was proven in [27] that when there are no recirculation zones in the flow, $T_{s}$ is equivalent to the third tortuosity considered in the present work
$T_{\Omega}=\frac{\displaystyle\int\limits_{\Omega}d\boldsymbol{x}\>V(\boldsymbol{x})}{\displaystyle\int\limits_{\Omega}d\boldsymbol{x}\>V_{0}(\boldsymbol{x})}=\frac{I[V]}{I[V_{0}]}.$ (8)
where the integration goes over the whole fluid volume $\Omega$ (see Fig. 6, rightmost plot). We will refer to this definition as the volume-integrated tortuosity. The notation $I_{\alpha}[\>\cdot\>]$ is a shorthand for the integral of some quantity over the fluid volume or its subset, i.e.
$I[f]_{\alpha}=\int\limits_{\Omega_{\alpha}}d\boldsymbol{x}\>f(\boldsymbol{x}),\quad\alpha\in\{\cdot,p,v\}$ (9)
and will be used further in the text. When the recirculation zones are present in the flow, the two latter tortuosities are not equal, rather the inequality holds [27]
$T_{s}\leq T_{\Omega}.$ (10)
Out of the two tortuosities $T_{s}$ and $T_{\Omega}$, the latter turns out to be significantly easier to calculate numerically. This is because of its locality – it does not make use of streamlines, which, due to numerical errors during their generation, can easily cross the boundary $\partial\Omega_{pv}$ between the percolating and recirculation volume during the integration along the fluid velocity field characteristics. This problem becomes especially pronounced in low-porosity porous media and at high Reynolds numbers, where a considerable portion of the fluid is occupied by the recirculation zones.
Figure 6: Schematic representation of the three definitions of tortuosities from Eqs. (5), (7), and (8). $T_{s}^{*}$ (left) and $T_{s}$ (middle) integrate over the surface elements $d\boldsymbol{x}$ of the intersections of an arbitrary surface AA and the corresponding fluid volumes. $T_{s}^{*}$ considers all the fluid volume $\Omega$, while $T_{s}$ only its percolating part, $\Omega_{p}$. $T_{\Omega}$ (right) integrates over the whole fluid volume $\Omega$. $T_{s}$ and $T_{\Omega}$ are used further in the present study.
2.3 Results analysis
2.3.1 Generation of streamlines
To generate the streamlines in SC samples, we numerically integrate the advection equation for the massless tracers advected in the fluid velocity field. We use the first-order explicit Euler method for the timestepping and meshless RBF-FD approximation [40] for the interpolation in space. Cubic radial functions augmented with second-order monomials are chosen as the interpolation basis, along with 25 closest neighbors in the approximation stencil. We use the C++ implementation of meshless approximation routines provided by Medusa library [41]. For the generation of streamlines, we use the normalized velocity field, $\boldsymbol{V}/V$, i.e., we set the magnitude of the velocity vectors to unity and preserve only their direction (see Fig. 7).
Figure 7: Schematic representation of the streamline generation procedure. The current increment of the streamline, $\Delta\boldsymbol{s}$ (denoted with an arrow), starts with the interpolation of the velocity to the current streamline end (blue filled circle). The interpolation uses a meshless scheme (and is thus naturally capable of operating on non-regular grids) and considers a number of the mesh points closest to the current streamline tip (filled squares, encircled in a dashed loop). The next streamline point (open circle) is calculated using the first-order Euler time integration.
The initial positions of the tracers depend on the task at hand. Due to the very different geometries and steady-state velocity fields in each of the SC samples, we tailor-cut the streamline generation for each case. For $\phi=0.1$ SC sample, for the calculation of the streamline-based tortuosity $T_{s}$, we place them uniformly at the fluid part of the $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers along $x_{1}$- and $x_{2}$-direction (see Fig. 8). To find the recirculation/percolating volumes, we place the tracers only in the vicinity of the solid boundaries at $x_{0}=0$ plane, with the distance $h_{s}=10^{-3}$ between the tracers. For this task, the grid from which the positions of the tracers are taken is rotated by $\pi/4$ about $x_{0}$-axis for a better compliance with the solid boundaries. For both tasks, we use the timestep length $\Delta t_{s}=10^{-3}$ during the integration along characteristics. Considering the initial positions of the streamlines for both tasks, theoretically, none of them should hit the solid wall, have a negative increment in $x$-direction, and all the streamlines should connect the inlet and the outlet (should be percolating). If any streamline fails to satisfy the mentioned conditions, we assume this is due to numerical errors and exclude it from further analysis.
Figure 8: Initial positions of the streamlines for the calculation of $T_{s}$ (left) and the identification of the recirculation/percolating volume (middle) for the low-porosity SC sample. The rightmost image shows which part of the domain is shown in the remaining ones. The fluid volume $\Omega$ is depicted in blue.
In case of $\phi=0.999$ SC sample, for the calculation of $T_{s}$ we place the tracers on a square grid of step $h_{s}=10^{-2}$ at the $x_{0}=0$ plane (see Fig. 9, top-left) and use the timestep length $\Delta t_{s}=10^{-2}$. We dismiss the streamlines that do not fulfill the previously stated criteria. Due to numerical errors in the solution of the governing equations, the mean values of $x_{1}$- and $x_{2}$-velocities in this case are not identically zero, although they are two orders of magnitude lower than the mean $x_{0}$-velocity. It does not impact the calculation of $T_{\Omega}$ significantly, however, it causes the nearly-straight streamlines to experience a drift in a direction perpendicular to $x_{0}$. This results in their elongation form $L$ by a certain value, the magnitude of which is comparable to the value of $T_{s}-1$. In turn, the streamline-based tortuosity is artificially increased to a non-negligible extent. Due to this, for the higher-porosity SC sample, in Eq. (6) we use the distance between the streamline’s start and end as the normalization length, rather than the porous sample’s side length $L$. For the determination of the recirculation/percolating volume, we use two sets of the initial positions of the tracers (see Fig. 9, top-right). To properly capture the flow separation on the rear part of the sphere’s surface, we place the tracers upwind from the sphere, in the vicinity of the domain’s centerline $x_{1}=x_{2}=0.5$ (see Fig. 9, bottom-left). After the integration, we apply the same criteria to dismiss certain streamlines as previously. To properly capture the saddle point at the end of the recirculation zone in the wake of the sphere, we place the tracers downwind from the sphere (also in the vicinity of the centerline) and integrate them in the upwind direction (see Fig. 9, bottom-right). We do not apply the previously described streamline exclusion criteria; rather,, we stop the integration before the streamlines reach the separation ring on the sphere’s surface, to avoid them falling into the recirculation zone due to the numerical errors. We use the separation $h_{s}=5\cdot 10^{-4}$ between the streamline starting positions along $x_{1}$- and $x_{2}$-direction and the timestep length $\Delta t_{s}=10^{-3}$ for the identification of the recirculation/percolating volume.
Figure 9: Initial positions of the streamlines for the calculations in the high-porosity SC sample. Top-left: for the calculation of $T_{s}$. Top-right: for the identification of recirculation/percolating volumes. Bottom row: magnified views at the initial positions of the forward- (A) and backward-integrated (B) streamlines. The $x_{1}=x_{2}=0.5$ line is shown for clarity.
2.3.2 Determining the recirculation and percolating volumes of fluid
To determine which points of the fluid domain $\Omega$ of the SC samples belong to the recirculation or percolating volume, the crucial part is to approximate the boundary between the two, $\partial\Omega_{pv}$. To do this, first we divide the domain into $N_{H}^{3}$ cubic cells and consider only those occupied by the fluid. We use $N_{H}=160$ for both systems. We then mark those of the cells which contain at least one point of the streamlines generated for the purpose of identification of $\Omega_{p}$ and $\Omega_{v}$ (Fig. 10, left). As mentioned above, we generate only the percolating streamlines (although not all of them). Due to this, holes may appear in the clusters of marked cells, denoted by an asterisk. Having obtained the cluster of the marked cells, we consider those lying on the cluster’s boundary (i.e. having at least one non-marked neighbor), near the expected position of $\partial\Omega_{pv}$, as the approximation of $\partial\Omega_{pv}$ (Fig. 10, middle). Lastly, all the cells lying on either side of the approximate boundary are marked as belonging to the recirculation or the percolating volume. The last step of this procedure automatically takes care of the proper coloring of the cells which do not contain any streamlines but should do so, if sufficiently many percolating streamlines were generated (the ones marked with an asterisk in Fig. 10). This way, one needs to generate the streamlines only in the vicinity of $\partial\Omega_{pv}$.
Figure 10: Schematic representation of the procedure for identifying the percolating ($\Omega_{p}$) and recirculation ($\Omega_{v}$) volumes. Each square represents a cell into which we divide the fluid volume. We note that those cells do not comply with the discretization. The percolating streamlines are plotted with thick dashed lines. We refer the reader to the paragraph Determining the recirculation and percolating volumes of fluid in Sec. 2.3 for the detailed explanation of the identification procedure.
2.3.3 Calculation of $T_{\Omega}$ constituents
To calculate the integrals of $V_{0}$ and $V$ over the whole volume of fluid in Eq. (8), in the case of stochastic porous samples, the first-order method, i.e., the summing of the values of the hydrodynamic fields from each lattice point, is used. In the case of SC samples, we use the ParaView utility IntegrateVariables [42, 43]. For the integration of $V_{0}$ and $V$ in the recirculation or the percolating volume in SC samples, we interpolate $V$ and $V_{0}$ to the centers of the cells used for the identification of $\Omega_{p}$ and $\Omega_{v}$. We use the same RBF-FD interpolation as in the generation of streamlines. Finally, we use the first-order method for the integration of the fields in each of the volumes.
3 Results
3.1 Difference in tortuosity–Reynolds number relation for porous samples of different porosities.
We first investigate the transition from the Darcy to inertial regime for two stochastic porous samples of porosities $\phi=0.7$ and $\phi=0.9$. Left subplot of Fig. 11 shows the dependence of the tortuosity $T_{\Omega}$ defined in Eq. (8) on the Reynolds number
$Re_{k}=\frac{\sqrt{k(Re_{k}\rightarrow 0)}\langle V_{0}\rangle\phi}{\nu},\quad\langle V_{0}\rangle=\frac{I[V_{0}]}{I[1]}$ (11)
where $k(Re_{k}\rightarrow 0)$ is the permeability at the vanishing Reynolds number. The same data are shown rescaled in the right subplot of Fig. 11, where by $T_{\Omega}(Re_{k}\rightarrow 0)$ we mean the value of $T_{\Omega}$ at the vanishing Reynolds number (denoted by the horizontal dashed line in the left subplot). We note that for $Re_{k}$ values beyond the investigated range, the solution went unsteady. We observe that in the Stokes regime, the tortuosity for each sample has a constant value, $1.07$ for $\phi=0.9$ and $1.175$ for $\phi=0.7$. As $Re_{k}$ increases, the tortuosity first decreases and then increases for both samples. However, the magnitude of the initial decrease and subsequent increase is considerably different for the two porosities. In the case of the higher-porosity sample, the difference between the minimum of $T_{\Omega}$ and the $T_{\Omega}(Re_{k}\rightarrow 0)$ is about $0.02$ ($0.3$ relative to the Darcy-regime value above unity, see the right subplot), and the increase is about $0.01$. For the lower-porosity sample, these values are $0.003$ ($0.01$ relative to the Darcy-regime value above unity) and $0.125$ ($0.67$ relative to the Darcy-regime value above unity), respectively. Such differences between the minimal and maximal values achieved by $T_{\Omega}$ in the steady-state regime were previously observed in stochastic high-porosity media [14] and in low-porosity geological samples [16]. The onset of inertial effects is also visible in the visualizations of the velocity fields, as the weaker channeling at the outlets of the samples (Fig. 12) and the emergence of the recirculation zones in the wakes of the obstacles (Fig. 13).
Figure 11: Left: the values of $T_{\Omega}$ for stochastic porous media of porosity $\phi=0.7$ (circles) and $\phi=0.9$ (squares) during the Darcy-inertial transition. The horizontal dashed lines denote the values of $T_{\Omega}$ in the limit of vanishingly small $Re_{k}$ ($Re_{k}\rightarrow 0$). Right: the same data as in the left subplot but rescaled by $T_{\Omega}$ at vanishingly small $Re_{k}$.
Figure 12: Streamlines of the velocity field for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is along $x_{0}$ axis. The shown streamlines are generated using the streamTracer utility of ParaView for the sake of the visualization only and are not used for the calculation of $T_{s}$ or the identification of percolating/recirculation volumes.
Figure 13: Velocity vectors at chosen sections for the stochastic porous media of porosity $\phi=0.7$ (top row) and $\phi=0.9$ (bottom row) at low (left column) and high (right column) $Re_{k}$. The flow is from left to right. For $\phi=0.7$, only every 3rd velocity vector in each direction is shown for clarity.
3.2 Splitting the volume-integrated tortuosity $T_{\Omega}$
To explain such a different behavior of $T_{\Omega}(Re_{k})$ in the low- and high-porosity stochastic systems, we first take a closer look at the definition of $T_{\Omega}$ and observe that both the numerator and denominator can be split into separate contributions coming from the percolating volume $\Omega_{p}$ and the recirculation volume $\Omega_{v}$
$T_{\Omega}=\frac{I[V]}{I[V_{0}]}=\frac{I[V]_{p}+I[V]_{v}}{I[V_{0}]_{p}+I[V_{0}]_{v}}.$ (12)
According to the definition of the vortex from Sec. 2, the mean velocity of the vortex in steady-state flow is zero, i.e., $I[\boldsymbol{V}]_{v}=\boldsymbol{0}$, so the term $I[V_{0}]_{v}$ in the denominator can be omitted. Thus, we can rewrite Eq. (12) as
$T_{\Omega}=\frac{I[V]_{p}}{I[V_{0}]_{p}}+\frac{I[V]_{v}}{I[V_{0}]_{p}}.$ (13)
Now, using the same procedure as in [27], it is possible to show that the surface integrals over $A\cap\Omega_{p}$ in Eq. (7) can be transformed into volume integrals over the entire percolating volume $\Omega_{p}$, i.e., $T_{s}=I[V]_{p}/I[V_{0}]_{p}$. This allows to rewrite Eq. (13) as
$T_{\Omega}=T_{s}+\frac{I[V]_{v}}{I[V_{0}]_{p}}.$ (14)
Such a form of the volume-integrated tortuosity allows one to clearly distinguish between the contributions from the percolating and the recirculation zones and directly points to the sources of inequality in Eq. (10). The denominator, $I[V_{0}]_{p}$, is proportional to the Reynolds number (cf. Eq. (11)) and does not introduce any new unknown variable into our considerations.
As noted previously, the proper identification of the percolating and recirculating volumes in the fluid is very sensitive to stray streamlines (i.e., those that, due to numerical errors, cross the boundary between the percolating and recirculating volumes). We observed that even several stray streamlines can significantly alter the value of the streamline-based tortuosity. Because of this, we conduct the numerical investigation of the two contributions to $T_{\Omega}$ on simple cubic (SC) systems, and use the obtained results to explain the behavior of $T_{\Omega}$ in the stochastic samples. SC samples can be thought of as models of particular fragments of a stochastic PM, at the same time having the geometry simple enough to allow for the accurate calculation of streamlines and identification of $\Omega_{p}$ and $\Omega_{v}$. To include in our analysis the cases with extremely high and low geometrical confinement, we consider SC systems of porosities $\phi=0.1$ and $\phi=0.999$. Exemplary visualizations of the steady-state velocity fields in the studied SC systems are shown in Figs. 14 and 15.
Figure 14: Visualizations of the sections of the velocity field in $\phi=0.1$ SC system for $Re_{k}=2.61\cdot 10^{-7}$ (left) and $Re_{k}=0.049$ (right). The gray surfaces are the obstacles’ walls, and only a few chosen walls are shown for clarity.
Figure 15: Visualizations of the sections of the velocity field in $\phi=0.999$ SC system for $Re_{k}=376$ (left) and $Re_{k}=1383$ (right). The gray surfaces are the obstacles’ walls.
3.3 Tortuosity as the inertia onset indicator in simple cubic systems
We continue our analysis by looking at the Reynolds number dependence of $T_{\Omega}$ and $T_{s}$ for the SC sample, shown in Fig. 16. In the Darcy regime ($Re_{k}\rightarrow 0$), the values of both tortuosities stay constant and are compliant with each other, as shown in [27]. We attribute small visible discrepancies between the two tortuosities to the errors introduced during the generation of streamlines and the identification of percolating/recirculation volumes. When the Reynolds number increases, the streamline-based tortuosity, $T_{s}$, decreases for both systems, which was observed in stochastic and periodic porous media by other authors [22, 15]. On the other hand, when the volume-based tortuosity, $T_{\Omega}$, is concerned, the inertial effects manifest themselves differently depending on the porosity, similarly to the stochastic porous media (cf. Fig. 11). In the lower-porosity case, its value decreases slightly in the range $Re_{k}\in[5\cdot 10^{-4};3\cdot 10^{-3}]$ to later increase significantly until the steady-state regime ends. Such behavior resembles the one observed for the lower-porosity stochastic porous media in Fig. 11 and also for a real-life porous sample [16]. For the higher-porosity sample, the volume-based tortuosity follows the streamline-based one closely and diverges from it only at $Re_{k}\approx 10^{3}$. Similarly to the lower-porosity system, the initial decline of $T_{s}$ is followed by its rise, to a much smaller extent, however. This is in line with the behavior of $T_{\Omega}$ for the higher-porosity stochastic porous media shown in Fig. 11, and in [14].
The strictly decreasing behavior of the streamline-based tortuosity can be attributed to the straightening of the percolating streamlines, which follows the growth of the recirculation zones volumes in the cavities and obstacles wakes. The volume-integrated tortuosity exhibits a more complex behavior, dependent on the sample’s porosity. Eq. (14) shows that the difference between $T_{\Omega}$ and $T_{s}$ is related to the integral of velocity magnitude inside the recirculation zones. Thus, we will further investigate the fall and the rise of the volume-based tortuosity in terms of the volume of the recirculation zones and the field of the velocity magnitude therein.
Figure 16: $T_{\Omega}$ and $T_{s}$ for $\phi=0.1$ and $\phi=0.999$ SC samples. The inset in the left subplot shows the magnification of the $\phi=0.1$ plot to the range of $T_{s}$ values. The horizontal dashed lines represent the values of $T_{\Omega}$ calculated for vanishingly small $Re_{k}$.
3.4 Origin of the $T_{\Omega}$ decrease and rise – volume of the recirculation zones
Let us first focus on the numerator of the recirculation volume-related term from Eq. (14), $I[V]_{v}/I[V_{0}]_{p}$. For convenience, we divide it by the total fluid volume, $I[1]$, i.e.
$\frac{I[V]_{v}}{I[1]}=\frac{I[1]_{v}}{I[1]}\cdot\frac{I[V]_{v}}{I[1]_{v}}$ (15)
where the first term on the right-hand side represents the vortex volume normalized by the total volume of fluid, and the other one – the total velocity magnitude per unit recirculation zone volume. We consider each of those two components separately. Figs. 17 and 18 show the visualization of the percolating volumes and recirculation zones for low- and high-porosity SC samples, respectively, for two different Reynolds numbers each. In both systems, as $Re_{k}$ rises, the recirculation zones grow in volume, which in the case of the low-porosity system is visible through the percolating volume shrinking to a straight jet connecting the inlet and the outlet. To quantify the changes in the size of the recirculation zones, we plot their volume, $I[1]_{v}$, normalized by the total volume of fluid, $I[1]$, in Fig. 19. In the low-porosity system, there already exist recirculation zones in the Darcy regime (in the channels perpendicular to the main flow, cf. left visualization in Fig. 14). The increase in their volume starts at $Re_{k}\approx 5\cdot 10^{-4}$, which corresponds to the start of $T_{\Omega}$ decrease in Fig. 16. The vortex volume then continues to grow until the end of the steady state regime, with significant slowdown for $Re_{k}>10^{-2}$, the value at which the slope of the $T_{\Omega}$ rise in Fig. 16 decreases. This reduction in the recirculation volume growth pace indicates the strong geometrical confinement for the vortex development. We highlight that at the highest $Re_{k}$, there already exist secondary vortices in the perpendicular channels (cf. right visualization in Fig. 14). The total increase in the volume of the recirculation zone is about 56% of the fluid volume, or – equivalently – $0.56\cdot 0.1/0.9\approx 0.062$ of the volume of the solid obstacles. On the other hand, in the high-porosity system, there is no recirculation in the Darcy regime (cf. left visualization in Fig. 15). The increase of its volume starts at $Re_{k}\approx 2\cdot 10^{2}$, which is about two decades later than the visible start of $T_{\Omega}$ decrease for this system in Fig. 16. This discrepancy can be related to the ‘formation of inertial cores’ [44] when the inertia already starts to change the velocity field, although no separation yet occurs. As the recirculation volume starts to grow, it does so with an increasing slope until the end of the steady-state regime (cf. right visualization in Fig. 15). The increase in the volume of the recirculation zone in this case is about $0.63$ of the solid obstacle volume. The growth of the recirculation zones in the case of high-porosity SC systems was highlighted by other authors [19] as the reason for lower drag compared to the flow over a single obstacle. Finally, we note that as the considered SC systems have a regular and simple geometry, phenomena like the non-strictly-monotonic growth of the volume of the recirculation zones [45] related to the instantaneous reorganization of the percolating flow paths are not visible here. Nevertheless, the start of the recirculation zones volume growth can still be related to the deviation from Darcy’s law at $Re\approx 10^{-3}$ (compare the data for $\phi=0.1$ SC system in Figs. 19 and 21). Interestingly, no high-wavenumber oscillations can be seen in the values of $T_{\Omega}(Re_{k})$ in the case of the stochastic samples (cf. Fig. 11), which suggests that the reorganization of the percolating flow paths might not impact the velocity integrals of the volume-integrated tortuosity.
Figure 17: Visualizations of the percolating volumes in $\phi=0.1$ SC system for the same values of $Re_{k}$ as in Fig. 14. The gray surfaces are the walls of the obstacles, and only a few chosen walls are shown for clarity. The visualizations use the histograms from which the percolating volumes were found. We show only the percolating volumes for clarity.
Figure 18: Visualizations of the recirculating volumes for $\phi=0.999$ SC system for the same values of $Re_{k}$ as in Fig. 15. The gray surfaces are the walls of the obstacles. The visualizations use the histograms from which the recirculation volumes were found. We show only the recirculation volumes for clarity.
Figure 19: Relative vortex volumes, $I[1]_{v}/I[1]$, for the two studied SC systems.
3.5 Origin of the $T_{\Omega}$ decrease and rise – energy confinement in the recirculation zones
The increase of the recirculation zone volume can explain the decrease of $T_{\Omega}$ because it directly implies the straightening of all the percolating streamlines and thus the decrease of $T_{s}$ in Eq. 14. However, it fails to address the rise of the volume-based tortuosity at the highest Reynolds numbers. This is why now we turn our attention to the integral of the velocity magnitude inside the vortex per vortex volume, $I[V]_{v}/I[1]_{v}$ (further in the text, we shall refer to this quantity as the specific velocity magnitude in vortex, for brevity). Its dependence on the Reynolds number is shown in Fig. 20 for both SC systems. We note that since there is no recirculation zone present in the higher-porosity SC system in the Darcy regime, we assume that the value of the specific velocity magnitude in the vortex is zero in those cases. For both SC systems, the specific velocity magnitude in vortices increases with Reynolds number. In the low-porosity sample, the rise proceeds from the lowest $Re_{k}$ with the first-order slope ($p=1$) and experiences a jump in the slope around $Re_{k}=3\cdot 10^{-3}$. It later attains the slope of the order approximately one again, in the range of $Re_{k}$ corresponding to those where the growth of $T_{\Omega}$ slows down (cf. Fig. 16). The initial first-order dependence is explained by the linearity of the Stokes equation and proportionality between the body force magnitude $g$ and the mean flow velocity $\langle\boldsymbol{V}\rangle$ in the Darcy regime (i.e., when $g$ is increased, the velocity field inside the vortex is only scaled proportionally by a constant). The subsequently higher slope indicates that a non-trivial mechanism involving the momentum transfer to the recirculation zone and the growth of its volume comes into play in the inertial regime. At the same time, an analogy between the Darcy regime and the inertial regime above $Re_{k}=10^{-2}$ can be drawn in the sense that both are characterized by the recirculation zone growing more in kinetic energy than volume. This further highlights the importance of the geometrical confinement in the emergence of inertial effects in low-porosity samples [16]. In the high-porosity system, the start of the rapid growth of the specific velocity magnitude in the vortex happens roughly at $Re_{k}$ when the recirculation zone volume starts to grow (cf. Fig. 19). It is interesting that even though the integral of the velocity magnitude over the whole fluid volume, $I[V]$, achieved at the highest Reynolds number is about $40$ times larger in the high-porosity system (results not shown in the plots), the specific velocity magnitude in vortex is about 2 times higher in low-porosity system ($\approx 140$ vs. $\approx 70$). This suggests that strong geometrical confinement leads to the accumulation of relatively large amounts of the kinetic energy in the recirculation zones, quantified also by other authors using the variance of the kinetic energy [13, 15, 14, 16].
In the end, in the low-porosity stochastic and SC systems, the greater kinetic energy confinement in the recirculation zones and the cessation of their volume growth at high Reynolds numbers lead to the more pronounced increase of $T_{\Omega}$, compared to the higher-porosity counterparts.
Figure 20: Integral of the absolute value of velocity in the vortex per vortex volume for SC systems (for brevity called the specific velocity magnitude in vortex in the text, $I[V]_{v}/I[1]_{v}$). The inset in the left subplot shows the same data but on a log-log scale.
4 Conclusions
The presented findings on the differences in behavior of the streamline-based and volume-integrated tortuosity during the Darcy-inertial transition in simple cubic porous media advance the understanding of the mechanisms behind the inertia onset in the studied porous systems. We show that porosity and the associated geometrical confinement are two of the key factors determining the behavior of the system in the inertial regime. The growth of the volume of the recirculation zones and the corresponding straightening of the percolating streamlines are captured by both types of tortuosity via their decrease in the initial stage of inertia onset. The later dynamics of the emergence of the inertial effects are strongly dependent on the system’s geometry. In low-porosity media, the geometrical confinement prevents the recirculation zones from growing freely in volume, thus mainly the kinetic energy confined within them increases. This is represented by the saturation of the streamline-based tortuosity at its minimal value, while the volume-integrated tortuosity grows rapidly in this regime. This is because the volume-integrated tortuosity contains information about the whole fluid volume, while the streamline-based one – only about the percolating volume. On the other hand, in high-porosity systems, the growth of the volume of the recirculation zones is not limited by the solid walls. Due to this, the increase of the volume-integrated tortuosity is less pronounced compared to the low-porosity samples, and the range of Reynolds numbers over which the two definitions of tortuosity are compliant with one another can extend almost to the end of the steady-state flow regime.
The investigated physical phenomena are closely related to the tools used to quantify them. The presented results show the utility of the two discussed measures of the inertial effects in porous media – streamline-based and volume-integrated tortuosity. Each of them separately carries information about various forms of the inertial effects in the flow – $T_{s}$ is more sensitive to the changes occurring solely in the percolating volume, while in $T_{\Omega}$ the contribution of the recirculation zones is also visible. The difference between the values of the two, as well as the magnitude of the growth of $T_{\Omega}$, can be treated as an indicator of the confinement of the recirculation zones. The presented results also highlight which aspects of the inertia onset in porous media are better captured by which definition of tortuosity and why, allowing for a more sensitive choice between the two in more application-oriented problems. One may stipulate, for instance, that the information about the recirculation zones carried by the volume-integrated tortuosity might prove useful when the mixing and reaction phenomena in the porous sample are of interest. On the other hand, the exclusive contribution of the percolating volume to the streamline-based tortuosity may carry more valuable information for studies concerned with the mass transport through the sample, e.g., during gas extraction, where the gaseous phase is often transported through preferential channels and the residuals of the fluid phase occupy the dead-end pores and cavities.
5 Acknowledgments
Funded by National Science Centre, Poland under the OPUS call in the Weave programme 2021/43/I/ST3/00228. This research was funded in whole or in part by National Science Centre (2021/43/I/ST3/00228). For the purpose of Open Access, the author has applied a CC-BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. D. Strzelczyk and M. Matyka acknowledge the financial support from the Slovenian Research And Innovation Agency (ARIS) research core funding No. P2-0095.
Appendix A Relation to other measures for the onset of inertia
To give a better perspective on our present findings, for the discussed stochastic and simple cubic systems, we additionally calculate other measures of the inertia onset. Fig. 21 shows the relation between the mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term (the Darcy’s law, Eq. (1)). A comparison with Figs. 11 and  16 shows that for the higher-porosity samples of each type the Reynolds number at which the mean velocity–forcing term relation becomes nonlinear corresponds roughly to the beginning of the fall of the tortuosities ($Re_{k}\approx 0.49$ and $5.75$ for the stochastic and simple cubic system, respectively). For the lower-porosity samples, the transition to the nonlinear dependence occurs later, around the minimum of $T_{\Omega}$ ($Re_{k}\approx 0.48$ and $1.45\cdot 10^{-3}$, for the stochastic and simple cubic system, respectively). Note also that for the simple cubic samples, the divergence from linearity in $\langle V_{0}\ranglegg$ relation is weaker than in the stochastic systems. This is especially in contrast to the behavior of $T_{\Omega}$ in the low-porosity SC sample, which experiences the biggest jump in value between the Darcy regime and the highest $Re_{k}$ from among all the studied systems. This suggests that the volume-integrated tortuosity might be a good indicator for the onset of inertia for extremely low-porosity systems.
Figure 21: Mean flow velocity in the $x_{0}$-direction versus the magnitude of the forcing term for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first-order scaling slopes anchored at the data points with the lowest value of $g$ (for some systems, those points are not visible at the presented axes ranges). The values at the arrows show the $Re_{k}$ for the data points around which the relation becomes non-linear. The acceleration and velocity for stochastic systems are given in lattice units.
Fig 22 shows the dependence of the inertial correction $f_{c,0}$ defined as [17]
$f_{c,0}=-\left(\frac{\langle\boldsymbol{V}\rangle}{|\langle\boldsymbol{V}\rangle|\phi}-\frac{k_{0}\boldsymbol{g}}{\mu|\langle\boldsymbol{V}\rangle|}\right)\cdot\hat{\boldsymbol{x}}_{0}=-1+\frac{k_{0}g}{\mu\langle V_{0}\rangle\phi}$ (16)
on the Reynolds number $Re_{k}$ for the studied stochastic and simple cubic systems. In Eq. (16), $\hat{\boldsymbol{x}}_{0}$ is the unit vector aligned with the $x_{0}$-axis, and the last equality holds under the assumption of the mean flow aligned with the $x_{0}$ direction, namely $\langle\boldsymbol{V}\rangle=\langle V_{0}\rangle\hat{\boldsymbol{x}}_{0}$. The weak inertia regime is characterized by the $f_{c,0}\propto Re_{k}^{2}$ scaling, while the strong inertia regime – by $f_{c,0}\propto Re_{k}$ scaling. In the case of stochastic samples, the latter regime is vaguely visible, which suggests that the transition between the two regimes takes place over the majority of the considered Reynolds numbers. In the case of simple cubic systems, the weak inertia regime is clearly visible for both porosities. In the several lowest $Re_{k}$ cases, the scaling becomes below-second order, which was also observed by other authors [17]. In the lower-porosity sample, the deviation from the weak inertia regime appears approximately at $Re_{k}=4\cdot 10^{-3}$, for which the $T_{\Omega}$ reaches its minimum (cf. Fig 16). For the highest Reynolds numbers, the relation changes into sublinear; however, it endures over a short range of $Re_{k}$ before the steady-state regime ends. It is possible that one observes the regime above strong inertia where $f_{c,0}\propto Re_{k}^{2}+Re_{k}$. One may relate this regime to the reduction in the increase of the slope of the volume-based tortuosity at $Re_{k}>2\cdot 10^{-2}$. For the higher-porosity simple cubic sample, the relation becomes sub-quadratic around $Re_{k}=5$, the value at which the tortuosities in Fig. 16 begin to rapidly decrease. Also in this case, for the highest Reynolds numbers, the $f_{c,0}(Re_{k})$ relation is clearly non-linear.
Figure 22: The inertial corection $f_{c,0}$ (Eq. (16)) versus the Reynolds number for the stochastic systems (left) and simple cubic systems (right). The dashed lines present the first- and second-order scaling slopes for the stochastic samples and 1/21/2– and second-order scaling slopes for the simple cubic samples. The inset in the right subplot shows the magnification of the highest-$Re_{k}$ region for the lower-porosity sample.
Appendix B Discretization independence study
To assess the discretization independence of $T_{\Omega}$ for the SC systems, we calculate its values from the solution of the governing equations obtained at different space discretization parameters ($h$ for uniform meshes and $h_{\text{min}}$ for refined meshes, cf. Section 2.1.3). For each porosity, the highest $Re_{k}$ is considered. In the case of the higher-porosity sample, we use the refined discretization and solve the transient problem. The results are presented in Table 1. The change of $T_{\Omega}$ between the subsequent discretizations is calculated relative to the value at the finer discretization.
$\phi=0.1$, $g=3.5\cdot 10^{6}$
$h$ $T_{\Omega}-1$ $\%$ change
$1/115$ $0.279$
$1/150$ $0.255$ $9.444$
$\mathbf{1/195}$ $\mathbf{0.296}$ $\mathbf{13.697}$
$1/253$ $0.299$ $0.996$
$1/330$ $0.301$ $0.822$
$\phi=0.999$, $g=5\cdot 10^{3}$
$h_{\text{min}}$ $\Delta t$ $T_{\Omega}-1$ $\%$ change
$1/116$ $2.25\cdot 10^{-6}$ $8.193\cdot 10^{-5}$
$\mathbf{1/150}$ $\mathbf{2.25\cdot 10^{-6}}$ $\mathbf{8.778\cdot 10^{-5}}$ $\mathbf{6.665}$
$1/196$ $2.25\cdot 10^{-6}$ $9.035\cdot 10^{-5}$ $2.851$
$1/254$ $1.5\cdot 10^{-6}$ $9.137\cdot 10^{-5}$ $1.112$
Table 1: Values above unity of the volume-integrated tortuosity $T_{\Omega}-1$ obtained for the highest forcing term $g$ for SC systems at various space discretization parameters used for the solution of the governing equations. The relative change in each row (rightmost column) is normalized by the value of $T_{\Omega}-1$ from the row above (one coarser discretization). The bold case row indicates the values discussed in the work.
We additionally perform the study of the independence of the $T_{s}$ value on the spacing $h_{s}$ and timestep value $\Delta t$ of the streamline generation procedure in SC samples. In this case, we take a look at the samples of the lowest $Re_{k}$, as these are the ones where the streamlines are expected to be the most tortuous, thus introducing larger errors at the segments of high curvature. We use the velocity fields obtained on the uniform $h=1/150$ or refined $h_{\text{min}}=1/150$ discretizations described in Sec. 2.1.3. We scale streamline seeding distance $h_{s}$ and the timestep for the integration along characteristics $\Delta t_{s}$ by a factor of $2$ between subsequent setups. The results are presented in Table 2
$\phi=0.1$, $Re_{k}=2.61\cdot 10^{-7}$
$h_{s}$ $\Delta t_{s}$ $T_{s}-1$ $\%$ change
 $4\cdot 10^{-3}$ $4\cdot 10^{-3}$ $0.0255$
$2\cdot 10^{-3}$ $2\cdot 10^{-3}$ $0.0244$ $4.598$
$\mathbf{10^{-3}}$ $\mathbf{10^{-3}}$ $\mathbf{0.0240}$ $\mathbf{1.676}$
$5\cdot 10^{-4}$ $5\cdot 10^{-4}$ $0.0238$ $0.861$
$\phi=0.999$, $Re_{k}=5.93\cdot 10^{-11}$
$h_{s}$ $\Delta t_{s}$ $T_{s}-1$ $\%$ change
$2\cdot 10^{-2}$ $2\cdot 10^{-2}$ $7.190\cdot 10^{-4}$
$\mathbf{10^{-2}}$ $\mathbf{10^{-2}}$ $\mathbf{7.343\cdot 10^{-4}}$ $\mathbf{2.085}$
$5\cdot 10^{-3}$ $5\cdot 10^{-3}$ $7.385\cdot 10^{-4}$ $0.561$
$2.5\cdot 10^{-3}$ $2.5\cdot 10^{-3}$ $7.404\cdot 10^{-4}$ $0.254$
Table 2: Values above unity of the volume-integrated tortuosity $T_{s}-1$ obtained for the lowest $Re_{k}$ for SC systems at various discretization parameters used for the streamline generation. The relative change in each row (rightmost column) is normalized by the value of $T_{\Omega}-1$ from the row above (one coarser discretization). The bold case row indicates the values discussed in the work.